# Inductive Meta-path Learning for Schema-complex Heterogeneous Information Networks

### v1

@@ 1 Introduction
Meta-paths are essential in analyzing HINs as they provide a higher-level description of the network structure that captures the relationships between different types of nodes and edges in the HINs. Given two entities, a meta-path refers to a sequence of entity types and the relation that exist between them. For example, in Fig. 1, the entity pair (Max Planck, Germany) is connected via following meta-paths:
@@
In this paper, we address aforementioned issues respectively by reinforcement learning (RL) and representation learning. The first issue of exploring the expansive meta-path space is tackled using RL, which offers an effective and flexible solution that can accommodate multiple training objectives for learning the meta-path discovery component. To achieve this, we parameterize the discovery component as an RL-based path-finding agent provided after the evaluation of each meta-path. This agent learns to selects plausible meta-paths by listening to rewards provided after meta-path evaluation.
@@
    We conduct extensive experiments, and demonstrate the superior ability of SchemaWalk in learning high-quality meta-paths for both KBs and general HINs. Impressively, SchemaWalk could effectively reason relations without finding specific evidence, with comparable performance to cases when these evidence for relations are provided.
@@ 2 Related Work
Multi-hop Reasoning. Multi-hop reasoning methods leverage the information stored in the path instances connecting entity pairs [21]. These methods mainly disregard relation types and operate on inference paths obtained by random walks [22, 23] or heuristics [24, 25]. Insighted by the advancement of deep RL, Xiong et al. firstly formulate the path-finding process as sequential decision making and describe an RL framework for learning multi-hop relational paths. Based on this, various papers explore different RL formulations and leverage different learning methods to boost multi-hop reasoning for prediction [20, 26].
Embedding-based Reasoning. Many KB completion tasks focus on learning vector representations for entities and relations with existing triples before using them for predictions. Most of these approaches ignore entity types and use different space for projection, e.g. TransE [27], KG2E [28] and RotatE [29]. However, these methods cannot capture compositional reasoning patterns and lack explanability [23]. Recent works have also explored the embedding of HINs, premised on discovered meta-paths. Dong et al. propose Metapath2Vec [30] to embed HINs with crafted meta-paths, but its performance are restricted by limited existing meta-paths. In particular, HIN2Vec [11] uses a multi-task learning approach to jointly embeds the information of different relations and instance graph structure.
@@ 3 Methodology
In this paper, we seek to learn meta-paths with high coverage and confidence. We thus start this section by introducing the definitions of HIN, meta-paths, and the most commonly-used evaluation metrics for measuring their plausibility including coverage and confidence. To enable SchemaWalk to perform inductive meta-path learning, we place the path-finding agent on the schema graph with trainable schema-level representations. Therefore, following the definitions, we introduce SchemaWalk in detail by focusing on the MDP formulation for meta-path learning and the agent design. Finally, we present the training algorithm.
@@ 3.1 Definitions and Objective
A meta-path is a path defined on the schema graph of a HIN. It serves as a vital tool in HIN analysis, capturing the semantic relationships between different types of entities in the network. A meta-path $M$ of length $l$ is defined as $M=t_{1}\xrightarrow{r_{1}}t_{2}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{l-1}}t_{l}$, where $t_{i}\in T$ denotes an entity type and $r_{i}\in R$ represents a relation. An instance-level path $P=v_{1}\xrightarrow{r_{1}}v_{2}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{l-1}}v_{l}$ satisfies $M$ if $\forall i\in\{1,\cdots,l\},t_{i}\in\tau(v_{i})$ and $\forall i\in\{1,\cdots,l-1\},r_{i}\in\phi(v_{i},v_{i+1})$. In this case, we also say $P$ is a path instance specific for $M$. The usage of meta-path instances is widespread for assessing the plausibility of meta-paths. Two commonly employed metrics for this purpose are coverage and confidence. Together, they provide a comprehensive measure of the usefulness of a meta-path.
@@
where $Con_{M}(v_{i},v_{j})$ indicates a meta-path instance that connects the entity pair $(v_{i},v_{j})$
@@
Given the entity pairs connected by a relation, coverage captures the frequency of a meta-path being satisfied by the paths connecting these entity pairs whilst confidence measures the correctness of a meta-path’s representation of the relation. A meta-path with high coverage and confidence is more likely to be relevant and useful. The complexity in computing these metrics for a meta-path involves finding all its path instances (also known as the evidence for it).
@@
Learning Objective. In this paper, we aim to learn meta-paths with high coverage and confidence for relations directly at the schema level. To achieve this goal, we train a path-finding agent on the schema graph to effectively navigate and explore the most promising meta-paths that lead to the relations. The evaluation of these meta-paths is performed using meta-path instances found in the instance graph. To demonstrate the effectiveness of our proposed methods across various scenarios, we assess the performance of our learned model using three distinct settings, as illustrated in Fig. 3. In the multi-relation inductive setting (Fig. 3 (A)), a single model is trained to discover meta-paths for multiple relations. The relations used in training is different from the test dataset, ensuring a disjoint relationship between them. Consequently, the meta-path instances for the relations in the test dataset are unseen during the training phase. Due to the inherent challenges in the multi-relation inductive setting, many existing methods struggle to effectively address it. To comprehensively evaluate our proposed methods, we include two additional settings that are comparatively less challenging. In the multi-relation transductive setting (Fig. 3 (B)), a single model is trained to discover meta-paths for multiple relations. The target relations used for training the model encompass the relations present in the test dataset. In the per-relation transductive setting (Fig. 3 (C)), a separate model is trained for each single relation to discover specific meta-paths exclusively tailored for that particular relation.
@@ 3.2 Reinforcement Learning Formulation
To fully unleash the expressive power of schema-level representations, we formulate the meta-path generator as a path-finding agent on the schema graph. Given the extensive meta-path space, we adopt an RL-based agent for efficient exploration. As depicted in Fig. 2 (Left), each time for the relation $r_{q}$, the agent randomly starts at a source type $t_{\rm src}$ in the entity type pairs $(t_{\rm src},t_{\rm tgt})$ connected by $r_{q}$ on the $\mathcal{G}_{S}$. This forms the query $r_{q}(t_{\rm src},t_{\rm tgt})=?$ for the agent who dynamically adjusts its policy to arrive at $t_{\rm tgt}$ to form meta-paths and simultaneously maximizing the rewards for the discovered meta-paths. We thus describe the meta-path learning problem, as a Markov Decision Process (MDP) on the schema graph by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$.
@@ 3.4 Training Pipeline
Due to the complex nature of meta-path finding, we leverage RL algorithm to train the policy network $\pi_{\theta}$ of the agent for $I_{base}$ iterations, on each relation for $I_{base}$ iterations, to maximize the rewards described in Eq. (4). To support multi-relation reasoning, we employ a periodic design that involves switching the relation for training every $I_{base}$ iterations, enabling each relation to undergo $I_{r}$ rounds of training. The training objective is,
@@
To estimate the gradient in Eq. (8), we randomly draw $K$ entity type pair $(t_{\rm src},t_{\rm tgt})$ connected by $r_{q}$ from $\mathcal{H}$ and run $N$ rollouts for each sample. The gradient is approximated by the sampled trajectories as follows,
@@
We append a moving average baseline to stabilize the training process by reducing the variance. The baseline is calculated by averaging the accumulative discounted rewards. Although the Actor-Critic pipeline with a parametric baseline prevails in the field of RL [33], we do not observe statistical significance by adopting it. Besides, to spur the exploratory behavior of SchemaWalk in discovering diverse meta-paths, we include an entropy regularization term with weight $\beta$ (with a decay rate) in the loss function. Ultimately, the ADAM optimizer is used to minimize the loss by rate $\alpha$.
@@ 4 Experiments
TABLE II: Meta-path space for some relations in YAGO26K-906
#Possible Metapaths #Valid Metapaths Valid Rate
wasBornIn $4,890,928$ $59,520$ $1.22\%$
LivesIn 3,937$,73$ 937 , 737 $29,751$ $0.76\%$
hasChild $2,172,857$ $12,469$ $0.57\%$
isCitizenOf $1,810,611$ $39,778$ $2.20\%$
PlaysFor $1,075,622$ $1,143$ $0.14\%$
isLeaderOf $1,006,833$ $9,511$ $0.94\%$
TABLE III: Inductive KG reasoning results on YAGO26K-906 and DB [^1] K-174, averaged over 5 runs. Enumeration cannot scale to DB [^1] K-174 and hence the results are not reported. The best/second best metrics are bolded/underline
@@ 4.1 Multi-relation Experiments
Datasets. We use two large KBs: YAGO26K-906 and Dbpedia for multi-relation settings. The meta-path space for some relations in YAGO26K-906 is showcased in Table II. We can see that, even in YAGO26K-906 which has comparatively fewer entity types and relations, each relation entails a million-scale meta-path space to search over and has extremely low valid rate. Here, a meta-path is valid if it has any meta-path instance in the $\mathcal{G}_{I}$, and low valid rate reflect the difficulty of mining meaningful meta-paths from the meta-path space. Besides, for DB [^1] K-174, which contains 305 relations, we train and test on sampled relations. A detailed description on the two datasets and the chosen train/test relations could be found at Appendix VII.
Inference. During inference, we utilize SchemaWalk to generate a set of high-quality meta-paths to reason about each test relation by employing a beam search with a large beam width of 400. Afterwards, the confidence of each outputted meta-path with regard to the queried relation $r_{q}$ is calculated, and is used as the score to infer tail entities. Given $(e_{h},r_{q},?)$, we apply the mined meta-paths on $e_{h}$ and rank possible tail entities after aggregating (max-pooling) the confidence of each meta-path. Entities that is inaccessible from $e_{h}$ via any meta-path are given a rank of infinite value. We use the standard metrics for KB completion tasks, i.e., the Hits@1, 3, 10 and the mean reciprocal rank (MRR). All results are averaged over five independent runs.
@@
SchemaWalk TransE DistMult ComplEx RotatE MINERVA RNNLogic
YAGO26K-906 Hits@1 0.157 0.092 0.039 0.069 0.180 0.133 0.161
Hits@3 0.250 0.225 0.052 0.123 0.260 0.214 0.263
Hits@10 0.377 0.316 0.118 0.174 0.347 0.310 0.364
MRR 0.223 0.176 0.059 0.106 0.237 0.192 0.225
DB [^1] K-174 Hits@1 0.668 0.538 0.359 0.626 0.768 0.684 0.671
Hits@3 0.845 0.809 0.481 0.728 0.822 0.786 0.754
Hits@10 0.875 0.858 0.555 0.769 0.849 0.863 0.783
MRR 0.754 0.675 0.431 0.682 0.799 0.745 0.716
@@ 4.1.1 Multi-relation Inductive Experiments
Observations. Table III displays the multi-relation inductive results on YAGO26K-906 and DB [^1] K-174. SchemaWalk significantly outperforms random walk in terms of efficiency and effectiveness. Even when we increase the search attempts for random walk by a factor of ten, it still cannot beat the accuracy of our approach on YAGO26K-906. On DB [^1] K-174, SchemaWalk exhibits more significant superiority over the random walk method. Enumeration present the best prediction answers on YAGO26K-906, but at an enormous computational cost (consuming nearly 498 times the inference time than SchemaWalk). Furthermore, enumeration cannot scale to DB [^1] K-174 given the enormous meta-path space. Notably, SchemaWalk produces meta-paths with significantly higher valid rate compared to other methods, which is about 6.5 times the valid rate obtained using random walk, and 13 times obtained using enumeration. This indicates that SchemaWalk indeed learns the mechanism to generate meaningful and relevant meta-paths.
@@ 4.1.2 Multi-relation Transductive Experiments
Baselines. We compare SchemaWalk with six widely-used KB reasoning baselines: TransE [27], DistMult [34], ComplEx [35], RotatE [29], MINERVA [20], RNNLogic [36]. RNNLogic could hardly scale to DB [^1] K-174 and we restrict the rule length to 3. Detailed information on their implementation could be referenced in Appendix A.
Observations. In Table IV, we show the multi-relation transductive results on YAGO26K-906 and DB [^1] K-174. SchemaWalk demonstrates superior performance over three embedding-based methods, namely TransE, DistMult, and ComplEx, and wins a noticeable advantage over the path-based MINERVA. On YAGO26K-906, SchemaWalk outperforms all other models in the Hits@10 metric and performs comparably with the rule-learning RNNLogic across all metrics. Additionally, on DB [^1] K-174, SchemaWalk achieves the highest scores for Hits@3, Hits@10, and the second-highest score for MRR. Overall, SchemaWalk competes strongly with state-of-the-art models for transductive KG query-answering.

### v2

@@ 1 Introduction
Meta-paths are essential in analyzing HINs as they provide a higher-level description of the network structure that captures the relationships between different types of nodes and edges in the HINs. Given two entities, a meta-path refers to a sequence of entity types and the relation that exists between them. For example, in Fig. 1, the entity pair (Max Planck, Germany) is connected via following meta-paths:
@@
In this paper, we address aforementioned issues respectively by reinforcement learning (RL) and representation learning. The first issue of exploring the expansive meta-path space is tackled using RL, which offers an effective and flexible solution that can accommodate multiple training objectives for learning the meta-path discovery component. To achieve this, we parameterize the discovery component as an RL-based path-finding agent provided after the evaluation of each meta-path. This agent learns to select plausible meta-paths by listening to rewards provided after meta-path evaluation.
@@
    We conduct extensive experiments, and demonstrate the superior ability of SchemaWalk in learning high-quality meta-paths for both KBs and general HINs. Impressively, SchemaWalk could effectively reason relations without finding specific evidence, with comparable performance to cases when these evidences for relations are provided.
@@ 2 Related Work
Multi-hop Reasoning. Multi-hop reasoning methods leverage the information stored in the path instances connecting entity pairs [21]. These methods mainly disregard relation types and operate on inference paths obtained by random walks [22, 23] or heuristics [24, 25]. Inspired by the advancement of deep RL, Xiong et al. firstly formulate the path-finding process as sequential decision making and describe an RL framework for learning multi-hop relational paths [26]. Based on this, various papers explore different RL formulations and leverage different learning methods to boost multi-hop reasoning for prediction [20, 27].
Embedding-based Reasoning. Many KB completion tasks focus on learning vector representations for entities and relations with existing triples before using them for predictions. Most of these approaches ignore entity types and use different space for projection, e.g. TransE [28], KG2E [29] and RotatE [30]. However, these methods cannot capture compositional reasoning patterns and lack explainability  [23]. Recent works have also explored the embedding of HINs, premised on discovered meta-paths. Dong et al. propose Metapath2Vec [31] to embed HINs with crafted meta-paths, but its performance is restricted by limited existing meta-paths. In particular, HIN2Vec [11] uses a multi-task learning approach to jointly embed the information of different relations and instance graph structure.
@@ 3 Methodology
In this paper, we seek to learn meta-paths with high coverage and confidence. We thus begin this section with an introduction to the definitions of HIN, meta-paths, and prevalent evaluation metrics, namely coverage and confidence, for assessing meta-path plausibility. To enable SchemaWalk to perform inductive meta-path learning, we place the path-finding agent on the schema graph with trainable schema-level representations. Therefore, following the definitions, we introduce SchemaWalk in detail by focusing on the MDP formulation for meta-path learning and the agent design. Finally, we present the training algorithm.
@@ 3.1 Definitions and Objective
A meta-path is a path defined on the schema graph of a HIN. It serves as a vital tool in HIN analysis, capturing the semantic relationships between different types of entities in the network. A meta-path $M$ of length $l$ is defined as $M=t_{1}\xrightarrow{r_{1}}t_{2}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{l-1}}t_{l}$, where $t_{i}\in T$ denotes an entity type and $r_{i}\in R$ represents a relation. An instance-level path $P=v_{1}\xrightarrow{r_{1}}v_{2}\xrightarrow{r_{2}}\cdots\xrightarrow{r_{l-1}}v_{l}$ satisfies $M$ if $\forall i\in\{1,\cdots,l\},t_{i}\in\tau(v_{i})$ and $\forall i\in\{1,\cdots,l-1\},r_{i}\in\phi(v_{i},v_{i+1})$. In this case, we also say $P$ is a path instance specific for $M$. The usage of meta-path instances is widespread for assessing the plausibility of meta-paths. In the field of association rule mining, coverage and confidence are two of the most crucial metrics for rule evaluation, each offering a distinct perspective [32]. Coverage denotes the rule’s applicability, while confidence serves as a gauge for its reliability. Here, we delve deeper these concepts and establish the rationale for combining these two metrics.
@@
where $Con_{M}(v_{i},v_{j})$ indicates a meta-path instance that connects the entity pair $(v_{i},v_{j})$.
@@
Given the entity pairs connected by a relation, coverage identifies the frequency of a meta-path being satisfied by the paths connecting these entity pairs whilst confidence measures the correctness of a meta-path’s representation of the relation. Meta-paths could display high accuracy but limited applicability (high confidence, low coverage), or broad applicability but low accuracy (high coverage, low confidence). By integrating these two metrics, we could identify rules that are of greater relevance and utility. The complexity in computing these metrics for a meta-path involves finding all its path instances (also known as the evidence for it).
@@
Learning Objective. In this paper, we aim to learn meta-paths with high coverage and confidence for relations directly at the schema level. To achieve this goal, we train a path-finding agent on the schema graph to effectively navigate and explore the most promising meta-paths that lead to the relations. The evaluation of these meta-paths is performed using meta-path instances found in the instance graph. To demonstrate the effectiveness of our proposed methods across various scenarios, we assess the performance of our learned model using three distinct settings, as illustrated in Fig. 3. In the multi-relation inductive setting (Fig. 3 (A)), a single model is trained to discover meta-paths for multiple relations. The relations used in training are different from the test dataset, ensuring a disjoint relationship between them. Consequently, the meta-path instances for the relations in the test dataset are unseen during the training phase. Due to the inherent challenges in the multi-relation inductive setting, many existing methods struggle to effectively address it. To comprehensively evaluate our proposed methods, we include two additional settings that are comparatively less challenging. In the multi-relation transductive setting (Fig. 3 (B)), a single model is trained to discover meta-paths for multiple relations. The target relations used for training the model encompass the relations present in the test dataset. In the per-relation transductive setting (Fig. 3 (C)), a separate model is trained for each single relation to discover specific meta-paths exclusively tailored for that particular relation.
@@ 3.2 Reinforcement Learning Formulation
As discussed in Section 2, current meta-path discovery methods adopt a bottom-top design. For example, MPDRL employs RL to navigate the instance graph, creating path instances prior to summarizing them as meta-paths. This approach not only faces challenges associated with the introduction of bias through path instances sampling but also prevents the direct learning of rules at the schema level. To fully unleash the expressive power of schema-level representations, we formulate the meta-path generator as a path-finding agent on the schema graph. Given the extensive meta-path space, we adopt an RL-based agent for efficient exploration. As depicted in Fig. 2 (Left), each time for the relation $r_{q}$, the agent randomly starts at a source type $t_{\rm src}$ in the entity type pairs $(t_{\rm src},t_{\rm tgt})$ connected by $r_{q}$ on the $\mathcal{G}_{S}$. This forms the query $r_{q}(t_{\rm src},t_{\rm tgt})=?$ for the agent who dynamically adjusts its policy to arrive at $t_{\rm tgt}$ to form meta-paths and simultaneously maximizing the rewards for the discovered meta-paths. We thus describe the meta-path learning problem, as a Markov Decision Process (MDP) on the schema graph by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$.
@@ 3.4 Training Pipeline
Due to the complex nature of meta-path finding, we leverage RL algorithm to train the policy network $\pi_{\theta}$ of the agent to maximize the rewards described in Eq. (4). To support multi-relation reasoning, we employ a periodic design that involves switching the relation for training every $I_{base}$ iterations, enabling each relation to undergo $I_{r}$ rounds of training. The training objective is,
@@
To estimate the gradient in Eq. (8), we randomly draw $K$ entity type pair $(t_{\rm src},t_{\rm tgt})$ connected by $r_{q}$ from $\mathcal{H}$ and run $N$ roll-outs for each sample. The gradient is approximated by the sampled trajectories as follows,
@@
We append a moving average baseline to stabilize the training process by reducing the variance. The baseline is calculated by averaging the accumulative discounted rewards. Although the Actor-Critic pipeline with a parametric baseline prevails in the field of RL [39], we do not observe statistical significance by adopting it. Besides, to spur the exploratory behavior of SchemaWalk in discovering diverse meta-paths, we include an entropy regularization term with weight $\beta$ (with a decay rate) in the loss function. Ultimately, the ADAM optimizer is used to minimize the loss by rate $\alpha$. With above training process, the policy network $\pi_{\theta}$ guides the learning of trainable representations of entity types and relations with the goal of maximizing the reward.
@@ 4 Experiments
MLN4KB
TABLE II: Inductive KG reasoning results on YAGO26K-906 and DB [^1] K-174, averaged over 5 runs. Enumeration cannot scale to DB [^1] K-174 and hence the results are not reported. The best/second best metrics are bolded/underline
@@
TABLE III: Meta-path space for some relations in YAGO26K-906
#Possible Metapaths #Valid Metapaths Valid Rate
wasBornIn $4,890,928$ $59,520$ $1.22\%$
LivesIn 3,937$,73$ 937 , 737 $29,751$ $0.76\%$
hasChild $2,172,857$ $12,469$ $0.57\%$
isCitizenOf $1,810,611$ $39,778$ $2.20\%$
PlaysFor $1,075,622$ $1,143$ $0.14\%$
isLeaderOf $1,006,833$ $9,511$ $0.94\%$
@@ 4.1 Multi-relation Experiments
Datasets. We use two large KBs: YAGO26K-906 and Dbpedia for multi-relation settings. The meta-path space for some relations in YAGO26K-906 is showcased in Table III. We can see that, even in YAGO26K-906 which has comparatively fewer entity types and relations, each relation entails a million-scale meta-path space to search over and has extremely low valid rate. Here, a meta-path is valid if it has any meta-path instance in the $\mathcal{G}_{I}$, and low valid rate reflects the difficulty of mining meaningful meta-paths from the meta-path space. Besides, for DB [^1] K-174, which contains 305 relations, we train and test on sampled relations. A detailed description on the two datasets and the chosen train/test relations could be found at Appendix VII.
Inference. During inference, we utilize SchemaWalk to generate a set of high-quality meta-paths to reason about each test relation by employing a beam search with a large beam width of 400. Afterwards, the confidence of each outputted meta-path with regard to the queried relation $r_{q}$ is calculated, and is used as the score to infer tail entities. Given $(e_{h},r_{q},?)$, we apply the mined meta-paths on $e_{h}$ and rank possible tail entities after aggregating (max-pooling) the confidence of each meta-path. Entities that are inaccessible from $e_{h}$ via any meta-path are given a rank of infinite value. We use the standard metrics for KB completion tasks, i.e., the Hits@1, 3, 10 and the mean reciprocal rank (MRR). All results are averaged over five independent runs.
@@
SchemaWalk TransE DistMult ComplEx RotatE MINERVA RNNLogic MLN4KB
YAGO26K-906 Hits@1 0.157 0.092 0.039 0.069 0.180 0.133 0.161 0.128
Hits@3 0.250 0.225 0.052 0.123 0.260 0.214 0.263 0.236
Hits@10 0.377 0.316 0.118 0.174 0.347 0.310 0.364 0.342
MRR 0.223 0.176 0.059 0.106 0.237 0.192 0.225 0.213
DB [^1] K-174 Hits@1 0.668 0.538 0.359 0.626 0.768 0.684 0.671 0.669
Hits@3 0.845 0.809 0.481 0.728 0.822 0.786 0.754 0.722
Hits@10 0.875 0.858 0.555 0.769 0.849 0.863 0.783 0.785
MRR 0.754 0.675 0.431 0.682 0.799 0.745 0.716 0.710
@@ 4.1.1 Multi-relation Inductive Experiments
Observations. Table II displays the multi-relation inductive results on YAGO26K-906 and DB [^1] K-174. SchemaWalk significantly outperforms random walk in terms of efficiency and effectiveness. Even when we increase the search attempts for random walk by a factor of ten, it still cannot beat the accuracy of our approach on YAGO26K-906. On DB [^1] K-174, SchemaWalk exhibits more significant superiority over the random walk method. Enumeration presents the best prediction answers on YAGO26K-906, but at an enormous computational cost (consuming nearly 498 times the inference time than SchemaWalk). Furthermore, enumeration cannot scale to DB [^1] K-174 given the enormous meta-path space. Notably, SchemaWalk produces meta-paths with significantly higher valid rate compared to other methods, which is about 6.5 times the valid rate obtained using random walk, and 13 times obtained using enumeration. This indicates that SchemaWalk indeed learns the mechanism to generate meaningful and relevant meta-paths.
@@ 4.1.2 Multi-relation Transductive Experiments
Baselines. We compare SchemaWalk with six widely-used KB reasoning baselines: TransE [28], DistMult [40], ComplEx [41], RotatE [30], MINERVA [20], RNNLogic [42] and MLN4KB [43]. RNNLogic could hardly scale to DB [^1] K-174 and we restrict the rule length to 3. Detailed information on their implementation could be referenced in Appendix A.
Observations. In Table IV, we show the multi-relation transductive results on YAGO26K-906 and DB [^1] K-174. SchemaWalk demonstrates superior performance over three embedding-based methods, namely TransE, DistMult, and ComplEx, and wins a noticeable advantage over the path-based MINERVA and the rule-based MLN4KB. On YAGO26K-906, SchemaWalk outperforms all other models in the Hits@10 metric and performs comparably with the rule-learning RNNLogic across all metrics. Additionally, on DB [^1] K-174, SchemaWalk achieves the highest scores for Hits@3, Hits@10, and the second-highest score for MRR. Overall, SchemaWalk competes strongly with state-of-the-art models for transductive KG query-answering.